Resistors for electronic circuits are manufactured on a high-speed automated machine. The machine is set up to produce a large run of resistors of 1,000 ohms each. To set up the machine and to create a control chart to be used throughout the run, 15 samples were taken with four resistors in each sample. The complete list of samples and their measured values are as follows SAMPLE NUMBER READINGS (IN OHMs) 985 1015 1005 1010 986 1003 1020 1015 993 005 1020 989 1001 992 979 006 101 003 984 981 989 14 013 1002 1005

Develop an X bar chart and an r-chart and plot the values. From the charts, what comments can you make about the process?
Answer #1
sample 1 2 3 4 Mean Range
1 1010 991 985 986 993 25
2 995 996 1009 994 998.5 15
3 990 1003 1015 1008 1004 25
4 1015 1020 1009 998 1010.5 22
5 1013 1019 1005 993 1007.5 26
6 994 1001 994 1005 998.5 11
7 989 992 982 1020 995.75 38
8 1001 986 996 996 994.75 15
9 1006 989 1005 1007 1001.75 18
10 992 1007 1006 979 996 28
11 996 1006 997 989 997 17
12 1019 996 991 1011 1004.25 28
13 981 991 989 1003 991 22
14 999 993 988 984 991 15
15 1013 1002 1005 992 1003 21
Average 999.1 21.733
X chart control limits N A2 D3 D4
UCL = x + A2R 2 1.88 0 3.27
LCL = x- A2R 3 1.02 0 2.57
4 0.73 0 2.28
5 0.58 0 2.11
R chart Control Limits 6 0.48 0 2
UCL = D4R 7 0.42 0.08 1.92
LCL = D3R 8 0.37 0.14 1.86
9 0.34 0.18 1.82
10 0.31 0.22 1.78
11 0.29 0.26 1.74
X BAR CHART
UCL = X + A2R = 999.1 + 0.73(21.733) = 1014.965
LCL = X - A2R = 999.1 - 0.73(21.733) = 989.235
R CHART
UCL = D4R = (2.28)(21.733) = 49.551
LCL = D3R = (0) (21.733) = 0
CONCLUSION
All the points are well within control limits
The process is controlled